# 1Huawei Technologies, 2NTU, 3HKU, 4CUHK

### v1

@@ 1 Introduction
Recent advances of large language models (LLMs) have led to remarkable progress in autonomous software engineering (SWE) agents, capable of repository-scale issue resolving (yang2024sweagent; zhang2024autocoderover; xia2024agentless). Unlike short, single-file tasks (du2023classeval; austin2021program; chen2021evaluating; li2022competition), SWE agents are able to navigate code repositories, faithfully reproduce failing tests, identify buggy files and lines, implement fixes, and validate within executable sandboxes (yang2023intercode; liu2024repobench; shrivastava2023repofusion), which demands long-horizon reasoning, robust tool use, and multi-turn interaction abilities.
To equip LLMs with these capabilities, a common approach is supervised fine-tuning (SFT), which distills knowledge from stronger SWE agents. However, prior efforts often lack datasets that include execution environments and realistic bug instances needed to produce high-quality, diverse trajectories (wang2025swe; yang2025swesmith; jain2025r2egym). Beyond SFT, performance on SWE tasks can be further improved via mid-training, but this typically requires substantially more computation and data (copet2025cwm; yang2025kimi). Alternatively, recent work applies reinforcement learning (RL) without a fixed teacher model or predefined SWE trajectories (deepswe2025; sonwane2025bugpilot; wei2025toward), yet limited executable instances remain a bottleneck. In addition, RL is more computationally intensive, demands heavier training infrastructure, and is prone to training collapse due to hyperparameter sensitivity. These practical constraints motivate a critical question: how far can we push a lightweight, SFT-only training recipe for SWE tasks?
@@
  • We introduce the SWE-Lego dataset, a collection of 32k high-quality task instances and 18k validated expert trajectories. The SWE-Lego dataset is constructed from two complementary sources: high-quality real-world GitHub pull requests (badertdinov2025swerebench), and scalable synthetic instances (yang2025swesmith), where both sources complement each other in quality and quantity: the real-world data provides the majority of benefits but is hard to scale, while the synthetic data provides additional gains with further scaling. Furthermore, we implement rigorous trajectory validation and filtering, such as preventing agents from Git hacking (i.e., inspecting Git history to recover solutions) to ensure the data teaches genuine problem-solving skills.
@@
  • We refine the conventional supervised fine-tuning to better train SWE agents. While standard SFT approaches have been widely applied for training (yang2025swesmith; wang2025swe), they typically treat every token in an expert trajectory equally during loss calculation. This overlooks a critical issue: trajectories often contain both successful actions and intermediate errors, and naively learning from mistakes could reinforce undesired behavior. To address this, we introduce step-level error masking, which excludes tokens associated with execution-time errors (e.g., tool failures or failing tests) from the loss. Moreover, we introduce another refinement of difficulty-based curriculum learning, which trains the model on easier tasks before progressing to harder ones, with difficulty estimated empirically by trajectory length.
@@
  • To trade additional compute for further gains, we study test-time scaling (TTS)  (jain2025r2egym; openhands_critic_model; deepswe2025) along two orthogonal dimensions: (i) sequential scaling by increasing the maximum number of interaction turns, and (ii) parallel scaling via multiple rollouts with verifier selection. We seek to explore the best TTS strategies for SWE tasks given the compute constraints. We find that sequential scaling saturates around 100–140 turns; beyond this point, compute is better allocated to parallel rollouts. In addition, generative verifiers (swe-gym; jain2025r2egym) consistently outperform regression-based scorers (openhands_critic_model) for parallel scaling.
@@
Empirically, SWE-Lego establishes new state-of-the-art results among open-source models on SWE-bench Verified (jimenez2024swebench): SWE-Lego-Qwen3-8B reaches 42.2%, and SWE-Lego-Qwen3-32B attains 52.6% using only SFT (Figure 1(a)). With test-time scaling, the performance rises to 49.6% and 58.8% under TTS@16, respectively. These results demonstrate that a meticulously crafted SFT pipeline can rival or exceed the performance of more complex and computationally expensive training regimes. Notably, our results are obtained without Git hacking, whereas prior reports are often inflated by such solution leakage. Ablations in Figure 1(b) show that our SWE-Lego dataset yields a substantial boost of +25.6%, refined SFT contributes an additional +3.8%, and TTS ultimately adds +6.2%, which together lift Qwen3-32B from 23.2% to 58.8%, confirming the effectiveness of each building blocks. We hope SWE-Lego can provide a reproducible and lightweight post-training paradigm for building effective SWE agents, and the associated data and models of this work will be open-sourced for future research in the community.
@@ 2.1 Overview
LLMs are increasingly integrated into autonomous agents for real-world SWE tasks. We adopt the SWE-Agent paradigm (yang2024sweagent), in which an LLM operates through a specialized agent-computer interface that allows it to autonomously navigate repositories, search and edit code, and execute tests to diagnose and resolve issues. In this context, a SWE task instance is usually defined as the concrete SWE issue within a codebase, comprising a natural-language problem description, test cases, a reference golden patch, and related artifacts. Besides, a trajectory is the complete sequence of actions taken by an agent and the interleaved observations from environments. To train capable SWE agents, our primary goal is to curate a large-scale, diverse, and executable collection of task instances, together with high-quality, validated trajectories.
@@ 2.2 SWE Task Instance Creation and Validation
The proposed dataset is built upon SWE-rebench (badertdinov2025swerebench), a large collection of real-world executable repositories. We select over 3,000 Python-centric repositories with permissive licenses and active maintenance histories. To ensure reproducibility, we deploy a fully automated pipeline that parses configuration files (e.g., setup.py) to build Docker containers. Only repositories that successfully build and pass sanity tests are retained, serving as the verifiable “base plates” for our subsequent task construction.
@@
Real-world tasks are derived from resolved GitHub pull requests (PRs) paired with corresponding issue descriptions. They offer high authenticity with production-level bug complexity, but are often labor-intensive and limited in quantity due to the strict filtering criteria (badertdinov2025swerebench; swe-gym; jain2025r2egym; zeng2025skywork). Moreover, each real-world task requires a unique sandbox aligned with the pre-fix snapshot. While SWE-smith (yang2025swesmith) proposes a “PR mirror” method to generate real-world instances (prompting LLMs to revert PR edits via .diff plaintext instead of checking out the base commits), we find this approach yields lower-quality data compared to directly collecting PRs that satisfy SWE-rebench (badertdinov2025swerebench)’s filtering rules. Thus, we adopt the latter strategy for real-world task construction.
Synthetic tasks, by contrast, are generated via active bug injection following SWE-smith (yang2025swesmith), leveraging two core techniques: (i) LLM Rewrite: prompting models to rewrite code using only function headers and docstrings; (ii) AST Reformulation: extracting abstract syntax trees (ASTs) for classes/functions and applying random transformations, e.g., removing conditionals/loops, modifying operators or dependencies. Synthetic tasks enable high scalability and efficiency, as they are not constrained by the availability of filtered real PRs. Even repositories with no qualified PRs can be used to create large-scale tasks at low cost. Additionally, synthetic tasks from the same repository share a single sandbox environment, since all bugs are injected into the same codebase version.
@@ 2.3 Trajectory Rollout and Validation
With the hybrid task instances, we now roll out expert trajectories that are later used for SFT. We adopt the OpenHands scaffold (v0.53) (wang2024openhands) with Qwen3-Coder-480B-A35B-Instruct (qwen3) as the teacher agent, and set the maximum number of interaction turns to 100. The model offers competitive coding and reasoning capabilities on challenging SWE tasks and, being open-weight, can be deployed locally at scale for trajectory distillation. During distillation, we found that the agent sometimes peeked at Git history or future commits, thereby cheating rather than faithfully resolving the issue. As showcased in Figure 4, we also observed avoidable tool-calling errors and tools that were ineffective in practice. To enhance trajectory quality, we propose and implement the following three practices.
@@
Pruning Ineffective Tools. Recent work (openai_gpt51_codex_max; anthropic_claude_sonnet_4_5) has begun integrating task-management tools such as task_tracker for long-horizon planning and progress tracking. However, we find that Qwen3-Coder-480B-A35B-Instruct struggles to use them effectively, which often cause execution errors (see Figure 4(c)). We therefore discard this tool and restrict the tool set to four essential operations: execute_bash, str_replace_editor, think, and finish, to keep trajectories streamlined, as evidenced by fewer interaction turns (from 65.5 to 62.5) and a higher valid-trajectory rate (from 27% to 30%) in Table 5.
@@ 3.2 Difficulty-Based Curriculum Learning
  • Model-Based Scoring: We fine-tune a Qwen2.5-Coder-32B-Instruct (hui2024qwen2) model as a difficulty scorer on a dataset of 1.5k human-annotated (task, difficulty) pairs. This scorer achieves 70% accuracy on a held-out test set and classifies tasks into three tiers: Easy (estimated solving time << 15 minutes), Medium (15 minutes to 1 hour), and Hard (>> 1 hour).
@@ 3.3 Training Details
For training, we use the Qwen3-8B/32B [^2] https://huggingface.co/Qwen/Qwen3-8B; https://huggingface.co/Qwen/Qwen3-32B (qwen3) and the internal Pangu-7B/32B series as the base models. We perform full-parameter SFT on the collected SWE-Lego trajectories using LLaMA-Factory framework (zheng2024llamafactory). Models are trained for 4 epochs with a global batch size of 64. We use the AdamW optimizer (loshchilov2019decoupledweightdecayregularization) with a weight decay of 0.01 and a cosine learning rate schedule with a warmup ratio of 0.1. The learning rate is set to 1e-4 for the 7B/8B models and 5e-5 for 32B models. To enable training with long-horizon trajectories, we use a maximum context length of 128k tokens. For the Qwen3 series, RoPE scaling techniques such as YaRN (peng2023yarnefficientcontextwindow) are adopted to support context lengths of up to 128k tokens.
@@ 4.2 Improved Parallel Scaling
During parallel scaling, a verifier takes a trajectory and its proposed patch and returns a score between 0 and 1 that reflects the likelihood of successful resolution. We compare two paradigms: the regressive verifier (openhands_critic_model), which appends a scoring head to the backbone and is trained with binary cross-entropy loss, and the generative verifier (swe-gym; jain2025r2egym), which formulates verification as text generation by predicting “yes” or “no” and computing the score from normalized token probabilities. The generative formulation aligns with the pre-trained next-token prediction objective, potentially better leveraging the model’s inherent knowledge (generative_verifiers). For each test instance, we generate $K$ candidate rollouts and let the verifier select the top-scoring one. TTS@$K$ is the fraction of instances resolved under this top-1 selection from $K$ candidates.
@@
We compare SWE-Lego-Verifier-8B against publicly available verifiers on SWE-Lego-Qwen3-8B rollouts (Figure 10(b)). Our verifier achieves TTS@16 of 49.6%, outperforming OpenHands-Critic-32B (44.0%) (openhands_critic_model) and R2E-Gym-Verifier-14B (47.0%) (jain2025r2egym). Beyond absolute performance, we observe qualitatively different scaling behaviors across verifier paradigms. OpenHands-Critic-32B, which employs the regressive paradigm, exhibits performance degradation at higher $K$, a counterintuitive result indicating that larger candidate pools overwhelm its discriminative capacity. In contrast, both generative verifiers (ours and R2E-Gym) maintain monotonic improvement toward the Pass@$K$ upper bound, further confirming that the generative formulation provides more robust scaling properties. The performance advantage over R2E-Gym-Verifier-14B likely stems from two factors: a larger volume of high-quality training trajectories and better distribution alignment between training data and SWE-Lego agent behaviors.
@@ 5 Comparisons with Existing Methods
We compare the performance of SWE-Lego models with both proprietary and open-source baselines on SWE-bench Verified. As mentioned in Section 2.2, Docker images of SWE-bench Verified released prior to Sep 3, 2025 may contain the ground truth commits, which can lead to Git hacking, e.g., via git log --all or by grepping commit messages. For fair comparisons with prior methods, we report our results both with and without git hacking. The results without hacking are obtained using the latest SWE-bench Verified Docker images. From Table 6, we see that our SWE-Lego-Qwen3-8B achieves a $42.2\%$ resolve rate with refined SFT and reaches 49.6% with further TTS@16, while SWE-Lego-Qwen3-32B attains $52.6\%$ with SFT and 58.8% with TTS@16. These hack-free results surpass most open-source models and several larger proprietary models, even when their results are generally inflated by git hacking. Our SWE-Lego-Pangu series also delivers competitive performance, further validating the effectiveness of our SWE-Lego dataset and approach.
@@
   OpenAI-GPT-4o (openai_gpt4o_2024) Internal pipeline - 33.2
   OpenAI-o3 (openai_o3_o4mini_2025) Mini-SWE-agent - 58.4
   Claude-4-Sonnet (anthropicClaudeSonnet) SWE-agent - 66.6
   Claude-4.5-Sonnet (anthropic_claude_sonnet_4_5) Mini-SWE-agent - 70.6
   OpenAI-GPT-5 (openai_gpt5_2025) OpenHands - 71.8
   Gemini 3 Pro Preview (google_gemini3_2025) Mini-SWE-agent - 74.2
@@
   Qwen2.5-Coder-Instruct-7B (hui2024qwen2) MOpenHands - 1.0
   Qwen3-8B (qwen3) OpenHands - 7.6
   SWE-Gym-7B (swe-gym) OpenHands SFT 10.6
   SWE-agent-LM-7B (yang2025swesmith) SWE-agent SFT 15.2
   Lingma-SWE-GPT-7B (ma2025swe) SWESynInfer SFT 18.2
   SWE-Mirror-LM-7B (wang2025swe) MOpenHands SFT 22.8
   SWE-Dev-7B (wang2025swed) OpenHands RL 23.4
   Klear-Agent-8B-SFT (mini-swe-agent-plus) Mini-SWE-agent-plus SFT 39.0
@@
   Qwen2.5-Coder-Instruct-32B (hui2024qwen2) MOpenHands - 6.2
   Qwen3-32B (qwen3) OpenHands - 23.2
   SWE-Gym-32B (swe-gym) OpenHands SFT 20.6
   R2E-Gym-32B (jain2025r2egym) R2E-Gym SFT 34.4
@@
   SWE-Dev-32B (wang2025swed) OpenHands RL 36.6
   Skywork-SWE-32B (zeng2025skywork) OpenHands SFT 38.0
@@
   SWE-agent-LM-32B (yang2025swesmith) SWE-agent SFT 40.2
   DeepSWE-32B-Preview (deepswe2025) OpenHands RL 42.2
@@
   SWE-Mirror-LM-32B (wang2025swe) MOpenHands SFT 52.2
   CWM-32B (copet2025cwm) Agentless CPT + SFT + RL 53.9

### v2

@@ 1 Introduction
Recent advances of large language models (LLMs) have led to remarkable progress in autonomous software engineering (SWE) agents, capable of repository-scale issue resolving (Yang et al., 2024; Zhang et al., 2024; Xia et al., 2024). Unlike short, single-file tasks (Du et al., 2023; Austin et al., 2021; Chen et al., 2021a; Li et al., 2022), SWE agents are able to navigate code repositories, faithfully reproduce failing tests, identify buggy files and lines, implement fixes, and validate within executable sandboxes (Yang et al., 2023; Liu et al., 2024; Shrivastava et al., 2023), which demands long-horizon reasoning, robust tool use, and multi-turn interaction abilities.
To equip LLMs with these capabilities, a common approach is supervised fine-tuning (SFT), which distills knowledge from stronger SWE agents. However, prior efforts often lack datasets that include execution environments and realistic bug instances needed to produce high-quality, diverse trajectories (Wang et al., 2025b; Yang et al., 2025c; Jain et al., 2025). Beyond SFT, performance on SWE tasks can be further improved via mid-training, but this typically requires substantially more computation and data (Copet et al., 2025; Yang et al., 2025d). Alternatively, recent work applies reinforcement learning (RL) without a fixed teacher model or predefined SWE trajectories (Luo et al., 2025; Sonwane et al., 2025; Wei et al., 2025), yet limited executable instances remain a bottleneck. In addition, RL is more computationally intensive, demands heavier training infrastructure, and is prone to training collapse due to hyperparameter sensitivity. These practical constraints motivate a critical question: how far can we push a lightweight, SFT-only training recipe for SWE tasks?
@@
  • We introduce the SWE-Lego dataset, a collection of 32k high-quality task instances and 18k validated expert trajectories. The SWE-Lego dataset is constructed from two complementary sources: high-quality real-world GitHub pull requests (Badertdinov et al., 2025), and scalable synthetic instances (Yang et al., 2025c), where both sources complement each other in quality and quantity: the real-world data provides the majority of benefits but is hard to scale, while the synthetic data provides additional gains with further scaling. Furthermore, we implement rigorous trajectory validation and filtering, such as preventing agents from Git hacking (i.e., inspecting Git history to recover solutions) to ensure the data teaches genuine problem-solving skills.
@@
  • We refine the conventional supervised fine-tuning to better train SWE agents. While standard SFT approaches have been widely applied for training (Yang et al., 2025c; Wang et al., 2025b), they typically treat every token in an expert trajectory equally during loss calculation. This overlooks a critical issue: trajectories often contain both successful actions and intermediate errors, and naively learning from mistakes could reinforce undesired behavior. To address this, we introduce step-level error masking, which excludes tokens associated with execution-time errors (e.g., tool failures or failing tests) from the loss. Moreover, we introduce another refinement of difficulty-based curriculum learning, which trains the model on easier tasks before progressing to harder ones, with difficulty estimated empirically by trajectory length.
@@
  • To trade additional compute for further gains, we study test-time scaling (TTS)  (Jain et al., 2025; OpenHands Team, 2025; Luo et al., 2025) along two orthogonal dimensions: (i) sequential scaling by increasing the maximum number of interaction turns, and (ii) parallel scaling via multiple rollouts with verifier selection. We seek to explore the best TTS strategies for SWE tasks given the compute constraints. We find that sequential scaling saturates around 100–140 turns; beyond this point, compute is better allocated to parallel rollouts. In addition, generative verifiers (Pan et al., 2025; Jain et al., 2025) consistently outperform regression-based scorers (OpenHands Team, 2025) for parallel scaling.
@@
Empirically, SWE-Lego establishes new state-of-the-art results among open-source models on SWE-bench Verified (Jimenez et al., 2024): SWE-Lego-Qwen3-8B reaches 42.2%, and SWE-Lego-Qwen3-32B attains 52.6% using only SFT (Figure 1(a)). With test-time scaling, the performance rises to 49.6% and 58.8% under TTS@16, respectively. These results demonstrate that a meticulously crafted SFT pipeline can rival or exceed the performance of more complex and computationally expensive training regimes. Notably, our results are obtained without Git hacking, whereas prior reports are often inflated by such solution leakage. Ablations in Figure 1(b) show that our SWE-Lego dataset yields a substantial boost of +25.6%, refined SFT contributes an additional +3.8%, and TTS ultimately adds +6.2%, which together lift Qwen3-32B from 23.2% to 58.8%, confirming the effectiveness of each building blocks. We hope SWE-Lego can provide a reproducible and lightweight post-training paradigm for building effective SWE agents, and the associated data and models of this work will be open-sourced for future research in the community.
@@ 2.1 Overview
LLMs are increasingly integrated into autonomous agents for real-world SWE tasks. We adopt the SWE-Agent paradigm (Yang et al., 2024), in which an LLM operates through a specialized agent-computer interface that allows it to autonomously navigate repositories, search and edit code, and execute tests to diagnose and resolve issues. In this context, a SWE task instance is usually defined as the concrete SWE issue within a codebase, comprising a natural-language problem description, test cases, a reference golden patch, and related artifacts. Besides, a trajectory is the complete sequence of actions taken by an agent and the interleaved observations from environments. To train capable SWE agents, our primary goal is to curate a large-scale, diverse, and executable collection of task instances, together with high-quality, validated trajectories.
@@ 2.2 SWE Task Instance Creation and Validation
The proposed dataset is built upon SWE-rebench (Badertdinov et al., 2025), a large collection of real-world executable repositories. We select over 3,000 Python-centric repositories with permissive licenses and active maintenance histories. To ensure reproducibility, we deploy a fully automated pipeline that parses configuration files (e.g., setup.py) to build Docker containers. Only repositories that successfully build and pass sanity tests are retained, serving as the verifiable “base plates” for our subsequent task construction.
@@
Real-world tasks are derived from resolved GitHub pull requests (PRs) paired with corresponding issue descriptions. They offer high authenticity with production-level bug complexity, but are often labor-intensive and limited in quantity due to the strict filtering criteria (Badertdinov et al., 2025; Pan et al., 2025; Jain et al., 2025; Zeng et al., 2025). Moreover, each real-world task requires a unique sandbox aligned with the pre-fix snapshot. While SWE-smith (Yang et al., 2025c) proposes a “PR mirror” method to generate real-world instances (prompting LLMs to revert PR edits via .diff plaintext instead of checking out the base commits), we find this approach yields lower-quality data compared to directly collecting PRs that satisfy SWE-rebench (Badertdinov et al., 2025)’s filtering rules. Thus, we adopt the latter strategy for real-world task construction.
Synthetic tasks, by contrast, are generated via active bug injection following SWE-smith (Yang et al., 2025c), leveraging two core techniques: (i) LLM Rewrite: prompting models to rewrite code using only function headers and docstrings; (ii) AST Reformulation: extracting abstract syntax trees (ASTs) for classes/functions and applying random transformations, e.g., removing conditionals/loops, modifying operators or dependencies. Synthetic tasks enable high scalability and efficiency, as they are not constrained by the availability of filtered real PRs. Even repositories with no qualified PRs can be used to create large-scale tasks at low cost. Additionally, synthetic tasks from the same repository share a single sandbox environment, since all bugs are injected into the same codebase version.
@@ 2.3 Trajectory Rollout and Validation
With the hybrid task instances, we now roll out expert trajectories that are later used for SFT. We adopt the OpenHands scaffold (v0.53) (Wang et al., 2024) with Qwen3-Coder-480B-A35B-Instruct (Yang et al., 2025a) as the teacher agent, and set the maximum number of interaction turns to 100. The model offers competitive coding and reasoning capabilities on challenging SWE tasks and, being open-weight, can be deployed locally at scale for trajectory distillation. During distillation, we found that the agent sometimes peeked at Git history or future commits, thereby cheating rather than faithfully resolving the issue. As showcased in Figure 4, we also observed avoidable tool-calling errors and tools that were ineffective in practice. To enhance trajectory quality, we propose and implement the following three practices.
@@
Pruning Ineffective Tools. Recent work (OpenAI, 2025a; Anthropic, 2025b) has begun integrating task-management tools such as task_tracker for long-horizon planning and progress tracking. However, we find that Qwen3-Coder-480B-A35B-Instruct struggles to use them effectively, which often cause execution errors (see Figure 4(c)). We therefore discard this tool and restrict the tool set to four essential operations: execute_bash, str_replace_editor, think, and finish, to keep trajectories streamlined, as evidenced by fewer interaction turns (from 65.5 to 62.5) and a higher valid-trajectory rate (from 27% to 30%) in Table 5.
@@ 3.2 Difficulty-Based Curriculum Learning
  • Model-Based Scoring: We fine-tune a Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) model as a difficulty scorer on a dataset of 1.5k human-annotated (task, difficulty) pairs. This scorer achieves 70% accuracy on a held-out test set and classifies tasks into three tiers: Easy (estimated solving time << 15 minutes), Medium (15 minutes to 1 hour), and Hard (>> 1 hour).
@@ 3.3 Training Details
For training, we use the Qwen3-8B/32B [^2] https://huggingface.co/Qwen/Qwen3-8B; https://huggingface.co/Qwen/Qwen3-32B (Yang et al., 2025a) series as the base models. We perform full-parameter SFT on the collected SWE-Lego trajectories using LLaMA-Factory framework (Zheng et al., 2024). Models are trained for 4 epochs with a global batch size of 64. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a weight decay of 0.01 and a cosine learning rate schedule with a warmup ratio of 0.1. The learning rate is set to 1e-4 for the 7B/8B models and 5e-5 for 32B models. To enable training with long-horizon trajectories, we use a maximum context length of 128k tokens. For the Qwen3 series, RoPE scaling techniques such as YaRN (Peng et al., 2023) are adopted to support context lengths of up to 128k tokens.
@@ 4.2 Improved Parallel Scaling
During parallel scaling, a verifier takes a trajectory and its proposed patch and returns a score between 0 and 1 that reflects the likelihood of successful resolution. We compare two paradigms: the regressive verifier (OpenHands Team, 2025), which appends a scoring head to the backbone and is trained with binary cross-entropy loss, and the generative verifier (Pan et al., 2025; Jain et al., 2025), which formulates verification as text generation by predicting “yes” or “no” and computing the score from normalized token probabilities. The generative formulation aligns with the pre-trained next-token prediction objective, potentially better leveraging the model’s inherent knowledge (Zhang et al., 2025). For each test instance, we generate $K$ candidate rollouts and let the verifier select the top-scoring one. TTS@$K$ is the fraction of instances resolved under this top-1 selection from $K$ candidates.
@@
We compare SWE-Lego-Verifier-8B against publicly available verifiers on SWE-Lego-Qwen3-8B rollouts (Figure 10(b)). Our verifier achieves TTS@16 of 49.6%, outperforming OpenHands-Critic-32B (44.0%) (OpenHands Team, 2025) and R2E-Gym-Verifier-14B (47.0%) (Jain et al., 2025). Beyond absolute performance, we observe qualitatively different scaling behaviors across verifier paradigms. OpenHands-Critic-32B, which employs the regressive paradigm, exhibits performance degradation at higher $K$, a counterintuitive result indicating that larger candidate pools overwhelm its discriminative capacity. In contrast, both generative verifiers (ours and R2E-Gym) maintain monotonic improvement toward the Pass@$K$ upper bound, further confirming that the generative formulation provides more robust scaling properties. The performance advantage over R2E-Gym-Verifier-14B likely stems from two factors: a larger volume of high-quality training trajectories and better distribution alignment between training data and SWE-Lego agent behaviors.
@@ 5 Comparisons with Existing Methods
We compare the performance of SWE-Lego models with both proprietary and open-source baselines on SWE-bench Verified. As mentioned in Section 2.2, Docker images of SWE-bench Verified released prior to Sep 3, 2025 may contain the ground truth commits, which can lead to Git hacking, e.g., via git log --all or by grepping commit messages. For fair comparisons with prior methods, we report our results both with and without git hacking. The results without hacking are obtained using the latest SWE-bench Verified Docker images. From Table 6, we see that our SWE-Lego-Qwen3-8B achieves a $42.2\%$ resolve rate with refined SFT and reaches 49.6% with further TTS@16, while SWE-Lego-Qwen3-32B attains $52.6\%$ with SFT and 58.8% with TTS@16. These hack-free results surpass most open-source models and several larger proprietary models, even when their results are generally inflated by git hacking.
@@
   OpenAI-GPT-4o (OpenAI, 2024) Internal pipeline - 33.2
   OpenAI-o3 (OpenAI, 2025c) Mini-SWE-agent - 58.4
   Claude-4-Sonnet (Anthropic, 2025a) SWE-agent - 66.6
   Claude-4.5-Sonnet (Anthropic, 2025b) Mini-SWE-agent - 70.6
   OpenAI-GPT-5 (OpenAI, 2025b) OpenHands - 71.8
   Gemini 3 Pro Preview (Google, 2025) Mini-SWE-agent - 74.2
@@
   Qwen2.5-Coder-Instruct-7B (Hui et al., 2024) MOpenHands - 1.0
   Qwen3-8B (Yang et al., 2025a) OpenHands - 7.6
   SWE-Gym-7B (Pan et al., 2025) OpenHands SFT 10.6
   SWE-agent-LM-7B (Yang et al., 2025c) SWE-agent SFT 15.2
   Lingma-SWE-GPT-7B (Ma et al., 2025) SWESynInfer SFT 18.2
   SWE-Mirror-LM-7B (Wang et al., 2025b) MOpenHands SFT 22.8
   SWE-Dev-7B (Wang et al., 2025a) OpenHands RL 23.4
   Klear-Agent-8B-SFT (Kwai-Klear, 2025) Mini-SWE-agent-plus SFT 39.0
@@
   Qwen2.5-Coder-Instruct-32B (Hui et al., 2024) MOpenHands - 6.2
   Qwen3-32B (Yang et al., 2025a) OpenHands - 23.2
   SWE-Gym-32B (Pan et al., 2025) OpenHands SFT 20.6
   R2E-Gym-32B (Jain et al., 2025) R2E-Gym SFT 34.4
@@
   SWE-Dev-32B (Wang et al., 2025a) OpenHands RL 36.6
   Skywork-SWE-32B (Zeng et al., 2025) OpenHands SFT 38.0
@@
   SWE-agent-LM-32B (Yang et al., 2025c) SWE-agent SFT 40.2
   DeepSWE-32B-Preview (Luo et al., 2025) OpenHands RL 42.2
@@
   SWE-Mirror-LM-32B (Wang et al., 2025b) MOpenHands SFT 52.2
   CWM-32B (Copet et al., 2025) Agentless CPT + SFT + RL 53.9